# Instruct, Not Assist: LLM-based Multi-Turn Planning and Hierarchical Questioning for Socratic Code Debugging

### v3

@@ Appendix A Evaluation Metrics
  • Average Number of Turns (Avg. Turns): We compute the average number of turns required by the method to reach the goal state. This metric provides insight into the efficiency and depth of the interaction process.
Table 3: Results on the 3-bug setting of the MULTI-DEBUG dataset compared with three ablation settings. TI indicates TreeInstruct, our method.
Syntactical Conceptual
Methods Avg. Turns Success Relevant Indirect Logic Success Relevant Indirect Logic
Vanilla [Uncaptioned image] 17.48 44.00 69.88 64.31 52.38 67.00 84.68 84.68 41.51
TI [Uncaptioned image]No Teaching 9.69 30.61 90.75 97.61 72.84 50.00 94.62 95.17 68.78
TI [Uncaptioned image]No State 16.34 25.51 51.61 97.21 41.09 53.00 47.57 94.70 20.36
Appendix B Human Expert Evaluators
@@
As mentioned before, our metrics in Tables 1 and 2 were obtained using volunteer human expert evaluators: two computer science teaching assistants with at least four years of high-school, undergraduate, and graduate-level teaching experience, with proficiency in Python and located in the USA. They were given the same set of instructions and the following set of guidelines:
@@
Below are some special cases/considerations the evaluators were also given:
@@
Appendix C Ablation Studies
Table 3 compares the results of the 3-bug setting with two ablation settings:
  • No Teaching: We remove the teaching functionality that kicks in when the Student has answered three consecutive questions incorrectly. Conversation is still guided by the state space representation, tree-based questioning, its updates, and the bug fixes proposed by the Student.
  • No State: We remove the state space representation. We guide conversation based on the conversation history, the previous questions asked, the Verifier feedback on the Student’s answer, and the bug fixes proposed by the Student.
Overall, we can see that when comparing the Llama-based ablations with the GPT4-based baselines, our ablation performance is still competitive, especially with respect to the relevance and indirectness of the questions. However, the significant drops in performance indicate the importance of our different modules, especially TI [Uncaptioned image]No State. We provide a detailed analysis of our ablation results below:
From TI [Uncaptioned image]to TI [Uncaptioned image]No Teaching, we see the Success rates and Logic scores to drop by 17.20% and 11.32%, respectively, on average. Teaching is a crucial part of our method because if the Student truly does not know a concept, then asking them more questions will not help them learn. While the Instructor’s job in TreeInstruct is to ask questions, at a certain point, it should teach the Student to clear their confusion. Hence, when we remove teaching, we see even fewer bugs solved and more repeated questions being asked, leading to low logical flow. However, we still see
Next, for TI [Uncaptioned image]No State, we see significant drops of 18.25% in Success rates, 46.63% in Relevance, and 51.39% in Logic scores, on average. The state space representation guides the question generation, ensures the questions are on topic to the bugs, and keeps track of the Student’s misunderstandings. Without this grounding, we noticed that the conversations (1) deviate from the real bugs– exploring areas such as time complexity optimization, which might not be the focus of the problem, (2) contain countless repeated questions that the Student already answered, and (3) jump from topic to topic abruptly in consecutive turns. These results show how impactful the state space representation is.
Appendix D Comparing State Space Representations in Multi-Bug Settings
@@
Tables 4, 5, and 6 outline the state space representations for the 1-bug, 2-bug, and 3-bug settings. As shown, 1-bug uses 3 states (states 1, 2, and 3) to solve the same but that 3-bug uses 1 state (state 1) to solve. This means the 1-bug state representation is much less compact than that for 3-bug.
@@
Table 7: Results on the side-by-side evaluation. Bolded and $\dagger$ values denote the top 2 comparisons respectively. Note: S-bug refers to the Socratic Debugging Benchmark. We abbreviate TreeInstruct as TI.
Comparison 3-bug 2-bug 1-bug S-bug
TI vs BRIDGE 71.43 68.00 83.67 94.63
TI vs Vanilla 100.00 90.00 69.39 50.33
BRIDGE vs Vanilla 57.14 62.00 40.82 24.83
Table 8: Results of human student evaluation across s(ingle)-bug (Socratic Debugging benchmark) and 3-bug (MULTI-DEBUG dataset) settings, broken down by the student level.
@@
Bugs Methods Avg. Turns Success Relevant Indirect Logic Success Relevant Indirect Logic
S-bug [Uncaptioned image] Level 1 6.0 100.00 66.67 66.67 100.00 100.00 91.67 100.00 50.79
Level 2 12.0 100.00 66.67 83.33 75.00 50.00 100.00 100.00 50.00
Level 3 8.0 0.00 87.50 100.00 50.00 100.00 67.50 90.00 42.50
Level 4 1.0 100.00 100.00 0.00 100.00 100.00 57.14 100.00 64.29
Level 5 1.0 100.00 100.00 100.00 100.00 100.00 100.00 75.00 75.00
3-bug [Uncaptioned image] Level 1 19.0 83.33 75.93 97.92 74.77 100.00 100.00 88.89 79.49
Level 2 11.7 83.33 100.00 100.00 78.57 100.00 100.00 86.67 82.50
Level 3 6.67 100.00 100.00 100.00 85.71 100.00 100.00 100.00 100.00
Level 4 4.7 100.00 93.33 100.00 76.67 100.00 100.00 83.33 88.89
Level 5 3.0 100.00 100.00 83.33 100.00 100.00 83.33 100.00 83.33
Appendix E Side by Side Evaluation
As mentioned in the main text, we perform a side-by-side evaluation to measure the percentage of times a user prefers our method TreeInstruct over the baselines baselines. Preference was measured as the average of all conversation metrics across syntactical and conceptual bugs. Based on the metrics, we assign each method a ranking (1, 2, or 3). Table 7 shows that TreeInstruct was preferred 68-94.6% of the time over the baselines. On average, TreeInstruct was preferred over BRIDGE 79.43% of the time, and over Vanilla 77.43% of the time.
Interpretation.
@@ Appendix F Interactive Evaluation with Human Students
  • Level 4: Student is new to TreeInstruct; they have been using Python for 2 years and is in their final year of undergraduate education in computer science.
  • Level 5: Student knows how TreeInstruct works; they act as an ally to intentionally provide good inputs so the method can resolve the bugs in as little turns as possible.
@@
    The students conversed with TreeInstruct until they were able to identify all of the bugs present in the code.
@@
Socratic questioning helped students learn programming concepts. The Level 3 student stated that, "If there was no conversation, I would be put off from attempting to fix and just try a bunch of different things based on the errors." Overall, students of Levels 2-4 (students with no knowledge of the system) were not able to identify all of the bugs before their interactions, but ended up solving them independently under the Socratic guidance of TreeInstruct.
Underlying model had a significant impact on user experience. Students had a significantly better experience with TreeInstruct [Uncaptioned image]compared to TreeInstruct [Uncaptioned image]. Specifically, the quality of the Verifier determined whether or not the questions posed by the Instructor would be overly repetitive or not.
@@ Appendix H License
All of the datasets used in this work, including our own, is under the Apache 2.0 License. Our use of existing artifact(s) is consistent with their intended use, specifically for the Socratic Debugging benchmark and in general, programming practice and feedback for the problems used in the MULTI-DEBUG dataset.
@@ Appendix I Prompts
The list should be ordered, with earlier attributes/tasks given priority over later ones (e.g., conceptual understanding tasks are a pre-requisite and thus more important than syntactical tasks). The following is an example of the state representation for the given example problem statement: example problem: Implement a Fibonacci sequence using recursion. {example buggy code} {example correct code} {example state representation}. Now do the same for the following problem statement, correct code, and student buggy code: {problem statement}, {correct code}, {buggy code}

### v4

@@ 4.4 Evaluation Metrics:
  • Average Number of Turns (#T): We compute the average number of turns required by the method to reach the goal state. This metric provides insight into the efficiency and depth of the interaction process.
4.5 Overall Results
In Tables 1 and 2, we see that with more structured representations of student knowledge and conversation state, TreeInstruct demonstrates significant improvements beyond the baselines. Across all multi-bug settings, we see an overall improvement of 16.6% and 11.59% in the success rates for syntactical and conceptual bugs, respectively. We also see an improvement of 13.47% and 14.89% for syntactical and conceptual bugs, respectively, across the three conversation metrics. For 1-bug, we see that the Vanilla [Uncaptioned image]baseline has the highest success for conceptual bugs, and the lowest Indirectness score, indicating that questions were very direct, and gave hints towards the bug fixes, which evidently increased the success rate. We see the same trend in the syntactical, single bug on Vanilla [Uncaptioned image]setting in Table 1. Overall, TreeInstruct demonstrates strong performance despite drastically different base models, [Uncaptioned image]and [Uncaptioned image]. TreeInstruct’s runtime performance is detailed in Appendix G.3.
Table 3: Results on the side-by-side evaluation. Bolded and $\dagger$ values denote the top 2 comparisons, respectively. Note: S-bug refers to the Socratic Debugging Benchmark. We abbreviate TreeInstruct as TI.
Comparison 3-bug 2-bug 1-bug S-bug
TI vs BRIDGE 71.43 68.00 83.67 94.63
TI vs Vanilla 100.00 90.00 69.39 50.33
BRIDGE vs Vanilla 57.14 62.00 40.82 24.83
Side-by-Side Evaluation:
Using the same evaluation metrics (Section 4.4), we performed a side-by-side evaluation that measures how often a user prefers TreeInstruct over the baselines. The results are shown in Table 3, in which we see that on average, TreeInstruct is given a higher ranking than BRIDGE 79.43% of the time, and a higher ranking than Vanilla 77.43% of the time. The details of this computation are in Appendix C. When specifically comparing TreeInstruct against BRIDGE, the key differences were that BRIDGE began with specific questions which revealed bug fixes and ended with general questions, lacking a logical flow in the conversation. In addition to being limited in a multi-bug setting, it posed less effective follow-up questions. In contrast, TreeInstruct prioritized high-level conceptual questions early on in order to build the foundation for later code-specific questions. Additionally, it asked for new information each turn and handled multiple bugs effectively. This is explicitly demonstrated in the example shown in Appendix B, Table 4.
Human Student Interaction:
We conducted a separate case study where human students directly interacted with TreeInstruct. We gathered five human volunteers of varying levels of programming backgrounds and knowledge, and presented each volunteer with three Single Bug problems and three 3-bug (MULTI-DEBUG) problems. The study shows that TreeInstruct can adapt to various levels of students effectively, as the scores are mostly comparable between our volunteers. The volunteers mentioned that TreeInstruct helped them learn programming concepts by forcing them to critically think about their mistakes, instead of trying random solutions. Also, our volunteers had better interactions with TreeInstruct [Uncaptioned image]than TreeInstruct [Uncaptioned image]because the [Uncaptioned image]Verifier is of higher quality. Appendix F includes further details of this study, with quantitative results in Table 10.
4.6 Ablation Studies
We perform two ablations on all 50 3-bug problems (bottom two rows of Table 2). For [Uncaptioned image]No State, we remove the state space representation, basing the question generation on the conversation history, Verifier feedback on the Student’s answer, and the Student-proposed bug fixes. For [Uncaptioned image]No Teaching, we remove the teaching functionality that kicks in after three consecutive incorrect Student answers. The conversation is still guided by the state space representation, tree-based questioning, its updates, and the Student-proposed bug fixes.
When compared to TreeInstruct [Uncaptioned image], No State success rates drop by 18.25%, Relevance by 46.63%, and Logic by 51.39%. Without the grounding provided by the state space representation, the conversations deviate from the real bugs and contain many repeated questions that the Student already answered. Additionally, when compared to TreeInstruct [Uncaptioned image], No Teaching success rates and Logic scores for drop by 17.20% and 11.32%, respectively. When the Student does not know some base-level of foundational knowledge, it is better to break the unnecessary cycle of asking questions and allow the Instructor to teach these concepts. Overall, these performance differences demonstrate the impact of each component.
Verifier Subtask Performance:
To understand the Verifier’s abilities to perform its designated subtasks (Section 3.1.2), given their complexity, we evaluate the Verifier on each subtask for the 3-bug setting. We demonstrate that for both [Uncaptioned image]and [Uncaptioned image], on the most difficult setting, both Verifier models feature high accuracy rates across all subtasks (86.3% average performance across all subtasks), most notably generateState, which shows that the base model does not have significant impact. Overall, this shows that TreeInstruct can reliably depend on the Verifier to complete such tasks. The detailed results are included in Appendix A.
4.6.1 Analysis
Fine-grained and dependency-aware state spaces improve conversation quality.
In our evaluation, we saw that state spaces typically contain 4–5 state variables, evenly distributed across all bugs. In other words, the number of bugs is inversely proportional to the number of variables allocated to each bug. Thus, 1-bug uses all state variables– each featuring a more fine-grained subtask to solve the bug, whereas 2- or 3-bug would use fewer variables for each bug (example provided in Appendix E). 1-bug’s fine-grained state variables lead to increased Success and Relevance scores as the conversation delves deep into the Student’s root misunderstanding of the bug. On the other hand, 3-bug features more coarse-grained state variables with a higher inter-bug dependency. Consider the (recursive) Fibonacci problem in Figure 1. Suppose the student is missing a base case and incorrectly calling the recursive function. Solving one bug requires adequate understanding of the other, thereby making it easier to solve. Another example can be found in Appendix B, Table 4. The 2-bug setting does not fully experience the benefits of fine-grained state variables or high bug-interdependency, resulting in a slight dip in scores: on average, a 7.62%, 0.83%, and 2.85% drop in Success, Relevant, and Logic scores, respectively– nonetheless, better than the baselines.
TreeInstruct tackles challenging problems more effectively.
Given the increase in problem difficulty between (1) the Socratic Debugging Benchmark (SDB) and MULTI-DEBUG (MD) datasets, and (2) from the 1 to 3-bug settings within MD, TreeInstruct is able to tackle challenging problems more effectively than baselines. The state space estimation breaks down problems into simple subtasks which the Instructor and Verifier can target using tree-guided questioning, even in complex conversations. Concretely, TreeInstruct’s performance in the SDB setting (Table 1) is comparable to that of the MD setting (Table 2), with only a 2.44% drop in performance from SDB to MD 1-bug and 1.53% drop from MD 1-bug to MD 3-bug. Comparatively, Vanilla [Uncaptioned image]drops by 14.55% from 1-bug to 3-bug. Overall, TreeInstruct’s Relevant, Indirect, and Logic scores remain high regardless of difficulty.
5 Conclusion
This paper proposes a novel method, TreeInstruct, for state space estimation and dynamic tree-based questioning for multi-turn Socratic instruction. We construct a novel multi-bug debugging dataset, MULTI-DEBUG, with 150 expert-annotated conceptual and syntactical problems and buggy solutions/fixes. Extensive experiments on an existing benchmark and MULTI-DEBUG demonstrate that TreeInstruct can be universally applied to both open and closed source models. We also showcase that TreeInstruct’s strong Socratic questioning abilities widely outperform all baselines through both (1) rigorous quantitative and qualitative expert evaluation (preferred over 77.94% of the time), and (2) real-world interactions with students of varying coding abilities.
6 Limitations & Future Work
Moreover, syntactical bugs might be “harder to see” for the language model, as it goes against the generation process to generate syntactically incorrect code. Breaking it down, a language model is trained to generate code with a colon at the end of for loops, if-statements, and method signatures, so if buggy code has a missing colon, the language model might ignore it. This results in syntactical bugs being harder to solve.
These limitations give way to exciting future work. Firstly, we can make use of vision language models to provide students with multi-modal teaching strategies, instead of relying solely on language. Additionally, we can enhance the framework, so it will explore new instruction methods when the questioning becomes cyclical. This can also help make the Instructor more reliable to generate consistent output across multiple runs on the same problem. Furthermore, we can utilize a structured fine-tuning approach to help the model better leverage the Verifier feedback and tree-based question generation process to make hierarchical Socratic planning and questioning inherent to a model. Overall, TreeInstruct can also be extended to automatically generalize to different teaching domains (e.g., quantitative reasoning).
7 Ethics Statement
8 Acknowledgements
Appendix A Breakdown of the Verifier Agent’s Subtask Performance
We perform ablation studies for each module on our most difficult setting (MULTI-DEBUG 3-bug), with both GPT-4 and Llama-3-8B Verifier Agents. For the sake of time, we used 20% of the dataset. For each subtask (generateState, verifyResponse, updateUnderstanding, isResolved), we define the following metrics of success:
  • generateState: the proportion of bugs that would be sufficiently resolved if the tasks (state space variables) were followed by the Student.
  • verifyResponse: a binary 0/1 for correct Verifier judgment and 0/1 for the correct explanation behind the Verifier’s judgment: $\frac{\text{verification\_score}+\text{explanation\_score}}{2}$.
  • updateUnderstanding: the proportion of state variables that were correctly validated (whether to update the state variable) and 0/1 for the correct explanation behind the Verifier’s judgment: $\frac{\text{validation\_score}+\text{explanation\_score}}{2}$.
  • isResolved: the proportion of Student proposed bug fixes which the Verifier correctly identifies as isomorphic/non-isomorphic to the corresponding ground truth bug fixes.
The results are in Table 5. Despite this being our most difficult setting, both Verifier models feature high accuracy rates across all subtasks, most notably generateState, which shows that the choice of base model does not have significant impact. Note: isResolved is primarily used for early stopping (e.g., if a student proposes all correct bugs during the first turn, then TreeInstruct stops early), so this subtask is not necessary. This subtask simply reduces computational costs.
Table 5: Performance of each Verifier subtask.
Subtasks Llama-3-8B GPT-4
generateState 90.0 100.0
verifyResponse 85.7 86.7
updateUnderstanding 88.6 85.9
isResolved 72.2 81.4
Appendix B Code Bug Interdependency
Table 4 shows an example of the effect of interdependent bugs, taken from real conversations that BRIDGE and TreeInstruct had with our student simulator Mistral-7B-Instruct. The questions aim to solve: (1) using a set instead of a list, and (2) using .add() instead of .append() to add items. As shown, Questions 1 and 2 tackle Bug 1, and Question 3 tackles Bug 2. The interdependency of changing the data structure first makes it easier to resolve the second bug, yielding higher success rates for the 3-bug setting.
Table 6: Inter-annotator agreement on the relevance, indirectness, logic scores, and bug type determination. Conducted on a random subset of the Single-Bug (SB) and MULTI-DEBUG 3-Bug (3-Bug) datasets.
Cohen’s $\kappa$ Agreement
Setting Success Relevant Indirect Logic Success Relevant Indirect Logic
SB 61.97 71.88 76.62 1.0 91.67 94.44 88.89 1.0
3-Bug 1.0 78.90 78.90 1.0 1.0 97.83 97.83 1.0
Appendix C Side by Side Evaluation
As mentioned in the main text, we perform a side-by-side evaluation to measure the percentage of times a user prefers our method TreeInstruct over the baselines. Preference was measured as the average of all conversation metrics across syntactical and conceptual bugs. Based on the metrics, we assign each method a ranking (1, 2, or 3). Table 3 shows that TreeInstruct was preferred 68-94.6% of the time over the baselines. On average, TreeInstruct was preferred over BRIDGE 79.43% of the time, and over Vanilla 77.43% of the time.
Interpretation.
Table 4 shows a comparison between the questions asked by BRIDGE and TreeInstruct. Appendix B specifies the details behind the targeted problem in the Table. On the left, we see that BRIDGE asks very similar questions regarding the design choice of using a list versus a set. TreeInstruct, on the right, asks one question regarding the design choice, one question about how to add elements to the correct data structure, and one intermediate question about why the design choice is suited towards the problem. Updating the state space, grounding questions towards the Student’s capabilities, and connecting concepts back to the problem are principles that rank TreeInstruct higher than BRIDGE.
Appendix D Human Expert Evaluators
@@
As mentioned before, the injected bugs into the dataset (Section 4.2) and evaluation in Tables 1 and 2 were obtained using volunteer human expert evaluators: two computer science teaching assistants with at least four years of high-school, undergraduate, and graduate-level teaching experience, with proficiency in Python and located in the USA. Their experiences equipped them with the first-hard experience of mistakes made by beginner programmers, leading to more realistic injected bugs. Furthermore, for evaluation, they were given the same set of instructions and the following set of guidelines:
@@
They were given a CLI-based annotation system. In our GitHub repository, we have provided the script (question_human_evaluation.py) and an example of the evaluation interaction (evaluation_example.pdf). Below are some special cases/considerations the evaluators were also given:
@@
On a random subset of the Single-Bug and MULTI-DEBUG 3-Bug datasets, we computed the inter-annotator agreement on the relevance, indirectness, and logic scores (Section 4.4) and bug type determination. We show these agreement results in Table 6.
Appendix E Comparing State Space Representations in Multi-Bug Settings
@@
Tables 7, 8, and 9 outline the state space representations for the 1-bug, 2-bug, and 3-bug settings. As shown, 1-bug uses 3 states (states 1, 2, and 3) to solve the same but that 3-bug uses 1 state (state 1) to solve. This means the 1-bug state representation is much less compact than that for 3-bug.
@@
Table 10: Results of human student evaluation across S(ingle)-bug (Socratic Debugging benchmark) and 3-bug (MULTI-DEBUG dataset) settings, broken down by the student level (Lvl).
@@
Bugs Lvl Avg. Turns Success Relevant Indirect Logic Success Relevant Indirect Logic
S-bug [Uncaptioned image] 1 6.0 100.00 66.67 66.67 100.00 100.00 91.67 100.00 50.79
2 12.0 100.00 66.67 83.33 75.00 50.00 100.00 100.00 50.00
3 8.0 0.00 87.50 100.00 50.00 100.00 67.50 90.00 42.50
4 1.0 100.00 100.00 0.00 100.00 100.00 57.14 100.00 64.29
5 1.0 100.00 100.00 100.00 100.00 100.00 100.00 75.00 75.00
3-bug [Uncaptioned image] 1 19.0 83.33 75.93 97.92 74.77 100.00 100.00 88.89 79.49
2 11.7 83.33 100.00 100.00 78.57 100.00 100.00 86.67 82.50
3 6.67 100.00 100.00 100.00 85.71 100.00 100.00 100.00 100.00
4 4.7 100.00 93.33 100.00 76.67 100.00 100.00 83.33 88.89
5 3.0 100.00 100.00 83.33 100.00 100.00 83.33 100.00 83.33
@@ Appendix F Interactive Evaluation with Human Students
  • Level 4: Student is new to TreeInstruct; they have been using Python for 2 years and are in their final year of undergraduate education in computer science.
  • Level 5: Student knows how TreeInstruct works; they act as an ally to intentionally provide good inputs, so the method can resolve the bugs in as little turns as possible.
@@
    The students conversed with TreeInstruct until they were able to identify all the bugs present in the code.
@@
Socratic questioning helped students learn programming concepts. The Level 3 student stated that, "If there was no conversation, I would be put off from attempting to fix and just try a bunch of different things based on the errors." Overall, students of Levels 2-4 (students with no knowledge of the system) were not able to identify all the bugs before their interactions, but ended up solving them independently under the Socratic guidance of TreeInstruct.
Underlying model had a significant impact on user experience. Students had a significantly better experience with TreeInstruct [Uncaptioned image]compared to TreeInstruct [Uncaptioned image]. Specifically, the quality of the Verifier determined whether the questions posed by the Instructor would be overly repetitive or not.
@@ G.2 Mistral and Llama
G.3 Runtime Comparison
We present the per-turn runtime for TreeInstruct (TI) across all multi-bug settings in Table 11. Llama was run on 1 NVIDIA Tesla V100 32GB, and OpenAI’s GPT-4 API was used for serverless inference. While the table indicates that Llama takes approximately 1 min per turn, this is due to a hardware bottleneck. With better hardware, TI-Llama should be able to run much faster to provide close to real-time tutoring. This is indicated by TI-GPT4 taking approximately 7 seconds per turn.
Table 11: Average runtime (in seconds) TreeInstruct takes for each turn, for each setting.
Setting Llama-3-8B GPT-4
1-bug 56.27 7.09
2-bug 65.69 7.77
3-bug 61.22 7.65
@@ Appendix H License
All the datasets used in this work, including our own, are under the Apache 2.0 License. Our use of existing artifact(s) is consistent with their intended use, specifically for the Socratic Debugging benchmark and in general, programming practice and feedback for the problems used in the MULTI-DEBUG dataset.
@@ Appendix I Prompts
The list should be ordered, with earlier attributes/tasks given priority over later ones (e.g., conceptual understanding tasks are a pre-requisite and thus more important than syntactical tasks). The following is an example of the state representation for the given example problem statement: example problem: Implement a Fibonacci sequence using recursion.
In-context example from Table 16
In the above example format (’example_explanation’ and ’example_state_representation’), output an ’explanation’ for your plan on which state_attributes to output and the ’state_representation’ with all NECESSARY ’state_attribute’s for the following ’problem’ statement based on ’correct_code’, the Student’s ’buggy_code’, the ’bug_description’s, and its respective ’bug_fixes’:
@@
Input:
example_problem:
Implement a fibonacci sequence using recursion.
example_buggy_code: def Fibonacci(n): if n <= 0: print("Incorrect input") elif n == 1: return 1 else: return Fibonacci(n-1)
example_bug_description:
On line 7, the function only recursively calls ‘Fibonacci(n-1)‘, which will then only return ‘1‘ from the edge case on line 5. Instead, the function should consider that the nth term of the Fibonacci sequence is computed as the sum of the preceding n-1 and n-2 values.
example_bug_fixes:
Replace ‘return Fibonacci(n-1)‘ with ‘return Fibonacci(n-1) + Fibonacci(n-2)‘ on line 7.
example_correct_code:
def Fibonacci(n): if n <= 0: print("Incorrect input") elif n == 1: return 1 else: return Fibonacci(n-1) + Fibonacci(n-2)
Output: example_explanation: Based on the ’bug_description’, the bug involves an incorrect recursive call, which indicates that the Student does not understand that the Fibonacci sequence requires taking the sum of the preceding two terms for getting the current value. An example where ‘n‘ is equal to ‘2‘ points this mistake out. Finally, based on the ’bug_fixes’, the Student must modify the recursive call to add ‘Fibonacci(n-2)‘ in order to reach the ‘correct_code‘ state.
example_state_representation:
state_attribute_1: Understand the definition of the Fibonnaci Sequence.
state_attribute_2: Consider the preceding two items in the sequence for computing the current value in the Fibonacci Sequence.
state_attribute_3: Consider the example when ‘n‘ is equal to ‘2‘.
state_attribute_4: Correctly recursively call ‘Fibonacci(n-2)‘ and add it to the existing ‘Fibonacci(n-1)‘.
Table 16: In context example for Table 15